# Supersymmetric BMS4 Algebras Revisited: Electric/Magnetic Superalgebras and Free Field Realization

### v1

@@ 3.3 Magnetic super BMS4 algebra
$\varphi^{(-\frac{3}{4},-\frac{3}{4})},\quad\phi^{(-\frac{1}{4},-\frac{1}{4})},\quad[\varphi^{(-\frac{3}{4},-\frac{3}{4})}_{r,\bar{r}},\phi^{(-\frac{1}{4},-\frac{1}{4})}_{s,\bar{s}}]=\frac{i}{4\pi^{2}}\delta_{r+s}\delta_{\bar{r}+\bar{s}},$ (2.9ch)
@@
$L_{m}=-4\pi^{2}i\sum_{s,\bar{s}}\left(s-\frac{1}{4}m\right)\varphi_{s,\bar{s}}\phi_{m-s,-\bar{s}}-\left(s+\frac{1}{4}m\right)\psi_{s,\bar{s}}\chi_{m-s,-\bar{s}}$ (2.9ci)
@@
$Q^{(\frac{1}{2},0)}_{r}=-4\pi^{2}i\sum_{s,\bar{s}}\varphi_{s,\bar{s}}\chi_{r-s,-\bar{s}}-\left(\frac{1}{2}s-\frac{3}{4}r\right)\phi_{s,\bar{s}}\psi_{r-s,-\bar{s}}$ (2.9cj)
@@
$R^{(1,1)}_{r,\bar{r}}=-4\pi^{2}i\sum_{s,\bar{s}}\phi_{s,\bar{s}}\bar{\phi}_{r-s,\bar{r}-\bar{s}}+\chi_{s,\bar{s}}\bar{\chi}_{r-s,\bar{r}-\bar{s}}.$ (2.9ck)
The transformation of these modes under the algebra is summarized in Table 3.3.
Table 7: Field modes transformation under Type II-II BMS4 algebra (2.9ah).
@@
$\mathcal{L}$ $=\frac{1}{2}\partial_{0}\phi\partial_{0}\phi-\psi_{+}\partial_{0}\psi_{-}-\psi_{-}\partial_{0}\psi_{+}-\bar{\psi}_{-}\partial_{0}\bar{\psi}_{+}-\bar{\psi}_{+}\partial_{0}\bar{\psi}_{-}+\bar{\varphi}\partial_{0}\bar{\phi}.$ (2.9cz)
@@
$\phi:l_{\phi}=0,\quad\Delta_{\phi}=\frac{1}{2},$ (2.9da)
@@
$\partial_{0}^{2}\phi=0,\quad\partial_{0}\psi_{+}=\partial_{0}\psi_{-}=0,\quad\partial_{0}\bar{\psi}_{+}=\partial_{0}\bar{\psi}_{-}=0,\quad\partial_{0}\bar{\varphi}=\partial_{0}\bar{\phi}=0,$ (2.9db)
@@
$\begin{aligned} &\bar{\varphi}(t,z,\bar{z})=\sum_{r,\bar{r}}\bar{\varphi}_{r,\bar{r}}z^{-r+\frac{1}{4}}\bar{z}^{-\bar{r}+\frac{1}{4}},\quad\bar{\phi}(t,z,\bar{z})=\sum_{r,\bar{r}}\bar{\phi}_{r,\bar{r}}z^{-r-\frac{5}{4}}\bar{z}^{-\bar{r}-\frac{5}{4}}.\end{aligned}$ (2.9df)
Realization of magnetic Type I-II algebra (2.9ai)
@@
$\varphi^{(-\frac{3}{4},-\frac{3}{4})},\quad\phi^{(-\frac{1}{4},-\frac{1}{4})},\quad[\varphi^{(-\frac{3}{4},-\frac{3}{4})}_{r,\bar{r}},\phi^{(-\frac{1}{4},-\frac{1}{4})}_{s,\bar{s}}]=\frac{i}{4\pi^{2}}\delta_{r+s}\delta_{\bar{r}+\bar{s}},$ (2.9dg)
@@
$L_{m}=-4\pi^{2}i\sum_{s,\bar{s}}\left(s-\frac{1}{4}m\right)\varphi_{s,\bar{s}}\phi_{m-s,-\bar{s}}-\left(s-\frac{3}{4}m\right)\psi_{s,\bar{s}}\chi_{m-s,-\bar{s}}$ (2.9dh)
@@
$Q^{(\frac{1}{2},0)}_{r}=-2\sqrt{2}\pi^{2}i\sum_{s,\bar{s}}2\left(\frac{1}{2}s-\frac{1}{4}r\right)\varphi_{s,\bar{s}}\chi_{r-s,-\bar{s}}+\phi_{s,\bar{s}}\psi_{r-s,-\bar{s}}$ (2.9di)
@@
$R^{(0,1)}_{r,\bar{r}}=2\sqrt{2}\pi^{2}i\sum_{s,\bar{s}}\phi_{s,\bar{s}}\bar{\phi}_{r-s,\bar{r}-\bar{s}}-\chi_{s,\bar{s}}\bar{\chi}_{r-s,\bar{r}-\bar{s}}.$ (2.9dj)
The modes transformation is shown in Table 3.3.
Table 8: Field modes transformation under Type I-II BMS4 algebra (2.9ai).
@@
$\mathcal{L}$ $=\frac{1}{2}\partial_{0}\phi\partial_{0}\phi-\psi_{+}\partial_{0}\psi_{-}-\psi_{-}\partial_{0}\psi_{+}+\psi_{-}\partial_{z}\psi_{-}-\bar{\psi}_{-}\partial_{0}\bar{\psi}_{+}-\bar{\psi}_{+}\partial_{0}\bar{\psi}_{-}+A_{+}\partial_{0}A_{-}.$ (2.9dy)
It is worth noting that this nonchiral symmetric theory includes spin-$\pm 1$ bosons $A_{\pm}$. Especially, the $A_{\pm}$ $\bar{\psi}_{\pm}$ part of this theory is single copy of the theory studied in [21] as perturbative supergravity. The spin and conformal dimensions of the fields are summarized as
$\phi:l_{\phi}=0,\quad\Delta_{\phi}=\frac{1}{2},$ (2.9dz)
@@
$\partial_{0}^{2}\phi=0,\quad\partial_{0}\psi_{+}=-\partial_{z}\psi_{-},\quad\partial_{0}\psi_{-}=0,$ (2.9ea)
@@
$\begin{aligned} &A_{+}(t,z,\bar{z})=\sum_{r,\bar{r}}\bar{\varphi}_{r,\bar{r}}z^{-r-\frac{3}{4}}\bar{z}^{-\bar{r}+\frac{1}{4}},\quad A_{-}(t,z,\bar{z})=\sum_{r,\bar{r}}\bar{\phi}_{r,\bar{r}}z^{-r-\frac{1}{4}}\bar{z}^{-\bar{r}-\frac{5}{4}}.\end{aligned}$ (2.9ee)
In the preceding constructions, we have realized the magnetic super BMS4 algebras through the theories (2.9at), (2.9be), (2.9ca), (2.9cz), and (2.9dy). Notably, there is no single simple theory that accommodates all types of magnetic algebras, highlighting their nontrivial structure compared to the electric algebras. Theories (2.9ca), (2.9cz), and (2.9dy) are not standard extended supersymmetric theories; nevertheless, the inclusion of $R$-symmetry is essential for the closure of the algebra. From this perspective, the left-hand and right-hand fermionic generators can be viewed as independent symmetry algebras, and their combinations should be regarded as extensions of the superalgebra. In each case, the $R$-symmetry maps bosons to bosons and fermions to fermions. It is also noteworthy that in (2.9ca), the presence of the Carrollian magnetic scalar demonstrates that, for supersymmetric BMS4 theories, the electric and magnetic sectors are equally significant and are related via $R$-symmetry. In the case of (2.9dy), the emergence of spin-$1$ fields corresponds to massless vector fields in the bulk 44D theory, rendering this nontrivial theory particularly relevant for future discussions in flat holography.
@@ 4 Discussion
In Section 3, we implemented an inverse construction to realize these superalgebras in free field theories. Notably, we identified a simple theory (2.9y) that realizes (2.9s), (2.9t), and (2.9u), consisting of one electric Carrollian scalar and two electric Carrollian spinors. For magnetic supersymmetry, there is a unique realization for each type. Although the of theory (2.9be) resembles the electric theory (2.9y), the conformal dimensions of the fermions differ, highlighting the distinct nature of magnetic constructions. The significance of magnetic Carrollian scalars is particularly evident in theory (2.9ca), while theory (2.9dy) plays a key role in the study of theories with bulk electromagnetic fields. Especially, part of (2.9dy) was studied in [21] as model of perturbative supergravity. The constructions presented in Section 3 focus on the simplest scalar supermultiplets, but the framework allows straightforward study of more general multiplets for further applications.
The global part of the BMS4 algebra is isomorphic to both the 44D Poincaré algebra and the 33D Carrollian conformal algebra. Similar relations hold for their supersymmetric extensions and the corresponding field theory realizations. As a supersymmetric extension of the 44D Poincaré algebra, it is imposed strict constraints by Haag–Łopuszański–Sohnius theorem. Indeed, the super Poincaré algebra coincides with the global part of the electric super BMS4 algebra (2.9s). The remaining electric superalgebras discussed in Section 2.2 involve supercharges with spin greater than $\tfrac{1}{2}$, and therefore their global parts do not correspond to physically realized superalgebras. Similarly, the magnetic superalgebras, whose supercharge anticommutators also produce superrotations, lack a direct physical superalgebra interpretation for their global sectors. Nevertheless, this does not diminish their relevance in the context of flat holography. On the contrary, these algebras highlight how supersymmetric extensions of BMS4 can go beyond the strict confines of super Poincaré, potentially capturing structures that only emerge in asymptotic or holographic regimes. In particular, such algebras may arise from large-radius limits of extended super asymptotic symmetries, and the appearance of $R$-symmetry offers a novel perspective on how internal symmetries interplay with asymptotic supersymmetry.
@@
Holography with supersymmetry. This paper has classified all possible super BMS4 algebras in the charge representation with finite subalgebras, providing a firm foundation for understanding supersymmetry in flat holography. These results serve as fundamental building blocks for constructing Fock spaces, boundary field theories, and analyzing constraints imposed by supersymmetry in holographic setups. In particular, the structure of $R$-symmetry offers guidance for constructing full 1010D spaces of string theories in flat holography, potentially shedding light on the embedding of flat space holography into string theory. Additionally, the results presented here suggest new directions for celestial holography: free superfield theories allow for explicit computation of correlation functions, thereby establishing bridges between supersymmetric scattering amplitudes and celestial correlators. Notably, the theory (2.9dy), which contains a massless spin-$1$ particle, is directly relevant for the study of soft photon theorems and may provide a testing ground for exploring how supersymmetry constrains celestial amplitudes.

### v2

@@ 3.3 Magnetic super BMS4 algebra
$\varphi^{(-\frac{3}{4},-\frac{3}{4})},\quad\phi^{(-\frac{1}{4},-\frac{1}{4})},\quad[\varphi^{(-\frac{3}{4},-\frac{3}{4})}_{r,\bar{r}},\phi^{(-\frac{1}{4},-\frac{1}{4})}_{s,\bar{s}}]=\frac{i}{4\pi^{2}}\delta_{r+s}\delta_{\bar{r}+\bar{s}},$ (3.54)
@@
$L_{m}=-4\pi^{2}i\sum_{s,\bar{s}}\left(s-\frac{1}{4}m\right)\varphi_{s,\bar{s}}\phi_{m-s,-\bar{s}}-\left(s+\frac{1}{4}m\right)\psi_{s,\bar{s}}\chi_{m-s,-\bar{s}}$ (3.55)
@@
$Q^{(\frac{1}{2},0)}_{r}=-4\pi^{2}i\sum_{s,\bar{s}}\varphi_{s,\bar{s}}\chi_{r-s,-\bar{s}}-\left(\frac{1}{2}s-\frac{3}{4}r\right)\phi_{s,\bar{s}}\psi_{r-s,-\bar{s}}$ (3.56)
@@
$R^{(1,1)}_{r,\bar{r}}=-4\pi^{2}i\sum_{s,\bar{s}}\phi_{s,\bar{s}}\bar{\phi}_{r-s,\bar{r}-\bar{s}}+\chi_{s,\bar{s}}\bar{\chi}_{r-s,\bar{r}-\bar{s}}.$ (3.57)
The transformation of these modes under the algebra is summarized in Table 7.
Table 7: Field modes transformation under Type II-II BMS4 algebra (2.34).
$\varphi_{s,\bar{s}}$ $\phi_{s,\bar{s}}$ $\psi_{s,\bar{s}}$ $\chi_{s,\bar{s}}$
$M_{r,\bar{r}}$ $\phi_{r+s,\bar{r}+\bar{s}}$ 0 0 0
$Q^{(\frac{1}{2},0)}_{r}$ $\left(-\frac{3}{4}r-\frac{1}{2}s\right)\psi_{r+s,\bar{s}}$ $\chi_{r+s,\bar{s}}$ $\varphi_{r+s,\bar{s}}$ $\left(-\frac{1}{4}r-\frac{1}{2}s\right)\phi_{r+s,\bar{s}}$
$Q^{(1,\frac{1}{2})}_{r,\bar{r}}$ $\chi_{r+s,\bar{r}+\bar{s}}$ 0 $\phi_{r+s,\bar{r}+\bar{s}}$ 0
$\bar{Q}^{(0,\frac{1}{2})}_{\bar{r}}$ $\left(-\frac{3}{4}\bar{r}-\frac{1}{2}\bar{s}\right)\bar{\psi}_{s,\bar{r}+\bar{s}}$ $\bar{\chi}_{s,\bar{r}+\bar{s}}$ $\left(-\frac{3}{4}\bar{r}-\frac{1}{2}\bar{s}\right)\bar{\varphi}_{s,\bar{r}+\bar{s}}$ $\bar{\phi}_{s,\bar{r}+\bar{s}}$
$\bar{Q}^{(\frac{1}{2},1)}_{r,\bar{r}}$ $\bar{\chi}_{r+s,\bar{r}+\bar{s}}$ 0 0 0
$R^{(1,1)}_{r,\bar{r}}$ $\bar{\phi}_{r+s,\bar{r}+\bar{s}}$ 0 $\bar{\chi}_{r+s,\bar{r}+\bar{s}}$ 0
@@
$\mathcal{L}$ $=\frac{1}{2}\partial_{0}\phi\partial_{0}\phi-\psi_{+}\partial_{0}\psi_{-}-\psi_{-}\partial_{0}\psi_{+}-\bar{\psi}_{-}\partial_{0}\bar{\psi}_{+}-\bar{\psi}_{+}\partial_{0}\bar{\psi}_{-}+\bar{\varphi}\partial_{0}\bar{\phi}.$ (3.58)
@@
$\phi:l_{\phi}=0,\quad\Delta_{\phi}=\frac{1}{2},$ (3.59)
@@
$\partial_{0}^{2}\phi=0,\quad\partial_{0}\psi_{+}=\partial_{0}\psi_{-}=0,\quad\partial_{0}\bar{\psi}_{+}=\partial_{0}\bar{\psi}_{-}=0,\quad\partial_{0}\bar{\varphi}=\partial_{0}\bar{\phi}=0,$ (3.60)
@@
$\begin{aligned} &\bar{\varphi}(t,z,\bar{z})=\sum_{r,\bar{r}}\bar{\varphi}_{r,\bar{r}}z^{-r+\frac{1}{4}}\bar{z}^{-\bar{r}+\frac{1}{4}},\quad\bar{\phi}(t,z,\bar{z})=\sum_{r,\bar{r}}\bar{\phi}_{r,\bar{r}}z^{-r-\frac{5}{4}}\bar{z}^{-\bar{r}-\frac{5}{4}}.\end{aligned}$ (3.64)
Realization of magnetic Type I-II algebra (2.35)
@@
$\varphi^{(-\frac{3}{4},-\frac{3}{4})},\quad\phi^{(-\frac{1}{4},-\frac{1}{4})},\quad[\varphi^{(-\frac{3}{4},-\frac{3}{4})}_{r,\bar{r}},\phi^{(-\frac{1}{4},-\frac{1}{4})}_{s,\bar{s}}]=\frac{i}{4\pi^{2}}\delta_{r+s}\delta_{\bar{r}+\bar{s}},$ (3.65)
@@
$L_{m}=-4\pi^{2}i\sum_{s,\bar{s}}\left(s-\frac{1}{4}m\right)\varphi_{s,\bar{s}}\phi_{m-s,-\bar{s}}-\left(s-\frac{3}{4}m\right)\psi_{s,\bar{s}}\chi_{m-s,-\bar{s}}$ (3.66)
@@
$Q^{(\frac{1}{2},0)}_{r}=-2\sqrt{2}\pi^{2}i\sum_{s,\bar{s}}2\left(\frac{1}{2}s-\frac{1}{4}r\right)\varphi_{s,\bar{s}}\chi_{r-s,-\bar{s}}+\phi_{s,\bar{s}}\psi_{r-s,-\bar{s}}$ (3.67)
@@
$R^{(0,1)}_{r,\bar{r}}=2\sqrt{2}\pi^{2}i\sum_{s,\bar{s}}\phi_{s,\bar{s}}\bar{\phi}_{r-s,\bar{r}-\bar{s}}-\chi_{s,\bar{s}}\bar{\chi}_{r-s,\bar{r}-\bar{s}}.$ (3.68)
The modes transformation is shown in Table 8.
Table 8: Field modes transformation under Type I-II BMS4 algebra (2.35).
$\varphi_{s,\bar{s}}$ $\phi_{s,\bar{s}}$ $\psi_{s,\bar{s}}$ $\chi_{s,\bar{s}}$
$M_{r,\bar{r}}$ $\phi_{r+s,\bar{r}+\bar{s}}$ 0 $2\left(-\frac{1}{4}r-\frac{1}{2}s\right)\chi_{r+s,\bar{s}}$ 0
$\bar{Q}^{(0,\frac{1}{2})}_{\bar{r}}$ $\left(-\frac{3}{4}\bar{r}-\frac{1}{2}\bar{s}\right)\bar{\psi}_{s,\bar{r}+\bar{s}}$ $\bar{\chi}_{s,\bar{r}+\bar{s}}$ $\left(-\frac{3}{4}\bar{r}-\frac{1}{2}\bar{s}\right)\bar{\varphi}_{s,\bar{r}+\bar{s}}$ $\bar{\phi}_{s,\bar{r}+\bar{s}}$
$\bar{Q}^{(\frac{1}{2},1)}_{r,\bar{r}}$ $\bar{\chi}_{r+s,\bar{r}+\bar{s}}$ 0 $2\left(-\frac{1}{4}r-\frac{1}{2}s\right)\phi_{r+s,\bar{r}+\bar{s}}$ 0
$R^{(0,1)}_{r,\bar{r}}$ $\frac{1}{\sqrt{2}}\bar{\phi}_{r+s,\bar{r}+\bar{s}}$ 0 $\frac{1}{\sqrt{2}}\bar{\chi}_{r+s,\bar{r}+\bar{s}}$ 0
@@
$\mathcal{L}$ $=\frac{1}{2}\partial_{0}\phi\partial_{0}\phi-\psi_{+}\partial_{0}\psi_{-}-\psi_{-}\partial_{0}\psi_{+}+\psi_{-}\partial_{z}\psi_{-}-\bar{\psi}_{-}\partial_{0}\bar{\psi}_{+}-\bar{\psi}_{+}\partial_{0}\bar{\psi}_{-}+A_{+}\partial_{0}A_{-}.$ (3.69)
It is worth noting that this nonchiral symmetric theory includes spin-$\pm 1$ bosons $A_{\pm}$. Especially, the $A_{\pm}$ and $\bar{\psi}_{\pm}$ part of this theory is single copy of the theory studied in [21] as perturbative supergravity. The spin and conformal dimensions of the fields are summarized as
$\phi:l_{\phi}=0,\quad\Delta_{\phi}=\frac{1}{2},$ (3.70)
@@
$\partial_{0}^{2}\phi=0,\quad\partial_{0}\psi_{+}=-\partial_{z}\psi_{-},\quad\partial_{0}\psi_{-}=0,$ (3.71)
@@
$\begin{aligned} &A_{+}(t,z,\bar{z})=\sum_{r,\bar{r}}\bar{\varphi}_{r,\bar{r}}z^{-r-\frac{3}{4}}\bar{z}^{-\bar{r}+\frac{1}{4}},\quad A_{-}(t,z,\bar{z})=\sum_{r,\bar{r}}\bar{\phi}_{r,\bar{r}}z^{-r-\frac{1}{4}}\bar{z}^{-\bar{r}-\frac{5}{4}}.\end{aligned}$ (3.75)
In the preceding constructions, we have realized the magnetic super BMS4 algebras through the theories (3.32), (3.39), (3.47), (3.58), and (3.69). Notably, there is no single simple theory that accommodates all types of magnetic algebras, highlighting their nontrivial structure compared to the electric algebras. Theories (3.47), (3.58), and (3.69) are not standard extended supersymmetric theories; nevertheless, the inclusion of $R$-symmetry is essential for the closure of the algebra. From this perspective, the left-hand and right-hand fermionic generators can be viewed as independent symmetry algebras, and their combinations should be regarded as extensions of the superalgebra. In each case, the $R$-symmetry maps bosons to bosons and fermions to fermions. It is also noteworthy that in (3.47), the presence of the Carrollian magnetic scalar demonstrates that, for supersymmetric BMS4 theories, the electric and magnetic sectors are equally significant and are related via $R$-symmetry. In the case of (3.69), the emergence of spin-$1$ fields corresponds to massless vector fields in the bulk 44D theory, rendering this nontrivial theory particularly relevant for future discussions in flat holography.
@@ 4 Discussion
In Section 3, we implemented an inverse construction to realize these superalgebras in free field theories. Notably, we identified a simple theory (3.21) that realizes (2.19), (2.20), and (2.21), consisting of one electric Carrollian scalar and two electric Carrollian spinors. For magnetic supersymmetry, there is a unique realization for each type. Although the of theory (3.39) resembles the electric theory (3.21), the conformal dimensions of the fermions differ, highlighting the distinct nature of magnetic constructions. The significance of magnetic Carrollian scalars is particularly evident in theory (3.47), while theory (3.69) plays a key role in the study of theories with bulk electromagnetic fields. Especially, part of (3.69) was studied in [21] as model of perturbative supergravity. The constructions presented in Section 3 focus on the simplest scalar supermultiplets, but the framework allows straightforward study of more general multiplets for further applications.
The global part of the BMS4 algebra is isomorphic to both the 44D Poincaré algebra and the 33D Carrollian conformal algebra. Similar relations hold for their supersymmetric extensions and the corresponding field theory realizations. As a supersymmetric extension of the 44D Poincaré algebra, it is imposed strict constraints by Haag–Łopuszański–Sohnius theorem. Indeed, the super Poincaré algebra coincides with the global part of the electric super BMS4 algebra (2.19). The remaining electric superalgebras discussed in Section 2.2 involve supercharges with spin greater than $\tfrac{1}{2}$, and therefore their global parts do not correspond to physically realized superalgebras. Similarly, the magnetic superalgebras, whose supercharge anticommutators also produce superrotations, lack a direct physical superalgebra interpretation for their global sectors. Nevertheless, this does not diminish their relevance in the context of flat holography. On the contrary, these algebras highlight how supersymmetric extensions of BMS4 can go beyond the strict confines of super Poincaré, potentially capturing structures that only emerge in asymptotic or holographic regimes. In particular, such algebras may arise from large-radius limits of extended super asymptotic symmetries, and the appearance of $R$-symmetry offers a novel perspective on how internal symmetries interplay with asymptotic supersymmetry.
@@
Holography with supersymmetry. This paper has classified all possible super BMS4 algebras in the charge representation with finite subalgebras, providing a firm foundation for understanding supersymmetry in flat holography. These results serve as fundamental building blocks for constructing Fock spaces, boundary field theories, and analyzing constraints imposed by supersymmetry in holographic setups. In particular, the structure of $R$-symmetry offers guidance for constructing full 1010D spaces of string theories in flat holography, potentially shedding light on the embedding of flat space holography into string theory. Additionally, the results presented here suggest new directions for celestial holography: free superfield theories allow for explicit computation of correlation functions, thereby establishing bridges between supersymmetric scattering amplitudes and celestial correlators. Notably, the theory (3.69), which contains a massless spin-$1$ particle, is directly relevant for the study of soft photon theorems and may provide a testing ground for exploring how supersymmetry constrains celestial amplitudes.